OpenStudy (anonymous):

how can you tell if an equation is an arithmetic series/sequence or a geometric series/sequence?

jimthompson5910 (jim_thompson5910):

a sequence is arithmetic if there is a common difference between each term example 1, 5, 9, 13, 17 each term has a common difference of 4 (1+5 = 4, 5+4 = 9, etc) notice how 5-1 = 4 9-5 = 4 13-9=4 17-13=4 so that's what they really mean when they say "common difference"

jimthompson5910 (jim_thompson5910):

with a geometric sequence, you have a common ratio example 7, 14, 28, 56 notice how 7*2 = 14 14*2 = 28 28*2 = 56 put another way 14/7 = 2 28/14 = 2 56/28 = 2 and we can see that the common ratio is 2

OpenStudy (anonymous):

Thanks you!

jimthompson5910 (jim_thompson5910):

np